I need to figure out how many possible milkshake combinations there can be with 64 ingredients. No ingredient can be used more than once in a combination and there is a max of 5 ingredients per combination.
Thanks for your help.
: I need to figure out how many possible milkshake combinations there can be with 64 ingredients. No ingredient can be used more than once in a combination and there is a max of 5 ingredients per combination.
Thanks for your help.

Answer by ankor@dixie-net.com(4684)

I need to figure out how many possible milkshake combinations there can be with 64 ingredients. No ingredient can be used more than once in a combination and there is a max of 5 ingredients per combination.
:
How about:
Five ingredients
64 * 63 * 62* 61 * 60 = 914,941,440
Four ingredients
64 * 63 * 62* 61 = 15,249,024
Three ingredients
64 * 63 * 62 = 249,984
Two ingredients
64 * 63 = 4032
1 ingredient
64
:
Add em up and you get: 930,444,544 combinations